OpenStudy (anonymous):

Divide 8 1/6 divided by 4 2/3. A 2 1/ 2 B 4/7 C 2 2/3 D 1 3/4

OpenStudy (anonymous):

start by converting the mixed numbers into improper fractions then remember, whenever you divide by a fraction, its the same as multiplying by the recipocal of that fraction ex.\[\frac{\frac{1}{2}}{\frac{1}{3}}=\frac{1}{2}*\frac{3}{1}=\frac{1*3}{2*1}=\frac{3}{2}\]